22. In a sample of 300 adults, 52 smoke. (22.a) Calculate the proportion of smokers in the sample. (Show your calculations.) (22.b) When we use Excel function =CONFIDENCE.T(α,σ,n), we need the value σ. Calculate that σ in this problem. (Show your calculations.) (22.c) Figure out the value E by using Excel function =CONFIDENCE.T(α,σ,n), with allowed error rate 0.05. (22.d) What is the 95% confidence interval of the proportion of smokers in population?
Answer #1

a)

proportion of smokers in the sample phat=52/300=0.1733

b)

σ =sqrt(p(1-p))=0.3785

c)

margin of error E= CONFIDENCE.T(0.05,0.3785,300) =0.043

d)

95% confidence interval of the proportion of smokers in population = phat -/+E =0.173-/+ 0.043 =0.130 to 0.216